OpenStudy (chillhill):

Solve (3^x)/(3^x+3)=1/3

OpenStudy (pawanyadav):

Multiply both side by (3^x+3) and try to solve

OpenStudy (pawanyadav):

Multiply both side by 3(3^x+3) You get 3^(x+1)=3^x+3 Now take log both side ( x+1)log3=xlog3+1